OpenStudy (anonymous):

given the distance between (x,1) and (-2,5) is 2 square root of 7. find the values of x. leave in simplest form

OpenStudy (anonymous):

(x+2)^2+(1-5)^2 = 4*7 ===> 28-16 = (x+2)^2 ===> x+2 =2sqrt(3) ==> x=2sqrt3-2

OpenStudy (anonymous):

that is correct but the other value from \[(x+2)^2=12\] \[x+2=\pm2\sqrt{3}\] \[x=-2\pm2\sqrt{3}\]i.e two answers are \[-2-2\sqrt{3}\] and \[-2+2\sqrt{3}\]

OpenStudy (anonymous):

thanks